OpenStudy (anonymous):

What is the value for x in 2(x + 3) = x + 6? If necessary, type No solution or Infinitely many solutions.

OpenStudy (anonymous):

use the distributive rule: 2(x+3) = 2x + 6 = x + 6, x=0